# Enabling Large-Scale and High-Precision Fluid Simulations on Near-Term Quantum Computers

### v2

@@ 3.3 Scaling quantum linear solver to arbitrary scale via the subspace method
To enhance scalability, we develop the subspace method inspired by the classical renowned Krylov subspace method and the generalized minimal residual (GMRES) method [41]. Instead of solving an $n$-dimensional linear system intractable large for near-term quantum computers, we iteratively build small spaces of lower dimension $m\ll n$ and find the approximate solution therein. More specific, we consider the $m$-order Krylov subspace $\mathcal{K}_{m}$ spanned by basis $\left\{\boldsymbol{b},A\boldsymbol{b},A^{2}\boldsymbol{b},...,A^{m}\boldsymbol{b}\right\}$. This integration of Iterative-QLS and subspace method is implemented via a quantum-classical hybrid procedure. As shown in Fig. 1, the subspace is constructed on classical computers, and a subspace linear system is generated and solved by quantum computers. The algorithm is presented in Alg. 2.
C provides theory details of the subspace method. We conduct convergence analysis and show that the algorithm can converge with a very loose error bound, i.e. $\epsilon<1$ for a single system solved by Iterative-QLS. Also, we show that the approximate solution $\hat{x}\in\mathcal{K}_{m}$ turns out to be a good estimation of $x$ when $m$ is large enough and with a restart scheme, detailed in C.2.
Iterative-QLS is chosen as the internal solver for the subspace method primarily due to its high-precision features. Replacing Iterative-QLS with a single VQLS process may cause divergence in the subspace process, as each integration is not guaranteed to sufficiently shrink the residual, considering the error produced in the real quantum computer. Furthermore, Iterative-QLS can be replaced by a faster and more accurate quantum solver, which remains a subject for future work.
The size of the subspace can be determined by the capability of the internal QLS process. By providing a more powerful quantum computer, one can utilize a larger subspace size to accelerate the iteration process, an idea widely discussed in the literature [42, 43, 44, 45, 46, 47]. Since QLS can solve $2^{n}$ linear system with $\operatorname{poly}(n)$ qubits, it offers a broad potential for quantum computing to accelerate at least one part of the fluid simulation. We conduct simulations on examples emulating scenarios with stronger quantum computers, detailed in F.
@@
1:Linear equations $A\boldsymbol{x}=\boldsymbol{b}$. Initial value $\tilde{\boldsymbol{x}}\neq A^{-1}\boldsymbol{b}$. Precision constraint $\epsilon$. Subspace dimension $m$.
2:Calculate residual $\boldsymbol{r}=\vec{b}-A\tilde{\boldsymbol{x}}$ and $\beta=\|\boldsymbol{r}\|_{2}$.
3:if  $\beta\leq\epsilon$ then
4:     Goto line 9.
5:end if
6:Construct subspace linear system $H\boldsymbol{y}=\boldsymbol{\xi}$ with coefficient matrix $V$ by GMRES.
7:Solve the subspace linear system through Iterative-QLS.
8:Update $\tilde{\boldsymbol{x}}=\tilde{\boldsymbol{x}}+V\boldsymbol{y}$.
9:Goto line 1.
@@ 3.4 Improving readout via sparse tomography
The concept of sparse sampling has been adopted in many domains which require fast conversion from quantum to classical data. $l_{\infty}$-norm tomography was the first proposed in constructing a deep convolutional neural network [40]. It offered an efficient method to generate such a sparse sample with $\mathcal{O}(\log N/\epsilon^{2})$ time complexity, where the error of each entry is less than $\epsilon$. It was further applied to quantum algorithms [5, 48, 49] that involves iterations where each iteration must convert quantum data to classical. In Ref. [5], the application of sparse tomography to various CFD problems are studied, showing that sparse tomography, acting as an “importance-sampling” scheme, is available for quantum-computing-based finite volume method.
However, $l_{\infty}$-tomography requires QRAM to store the output and perform state preparation, as well as an additional qubit, resulting in extra costs for implementing the controlled version of the entire process. QRAM is unrealistic in the near term, thus $l_{\infty}$-tomography is impracticable in near-term quantum computers. To address these limitations, a new method called “sparse tomography” is introduced in this study to achieve the same sparse sampling task without utilizing QRAM or any additional qubits.
The major idea of sparse tomography is the exploitation of the shadow tomography [50, 51], which is a novel tool to measure $O(\log N)$ time to compute $N$ expectation values by exploiting a classical postprocessing procedure named “classical shadow”. In our algorithm, we construct the classical shadow by defining a computation graph to determine the sign of the quantum state. The algorithm’s procedure is shown in Alg. 3 and the complexity and proof can be found in D. Because the algorithm does not use QRAM and it only involves measurements on various measurement bases (refer to Ref. [51] for the concept about measurement bases), this procedure is efficient and completely capable for NISQ devices.
@@
1:A quantum algorithm that produces a real-valued state $|x\rangle$.
2:Sample $|x\rangle$ with $36\log N/\epsilon^{2}$ copies to produce a sparse probability vector $\boldsymbol{p}$, and the position vector $\boldsymbol{v}$
3:Construct graph $\mathcal{G}=(V,E)$ with $k=\log\delta/\log v(1-p)$ random connected spanning subgraphs, where each subgraph contains all nodes with least edges ($d-1$ edges). For each subgraph, prepare all sign-determination Hamiltonians $\{H_{X}^{i,j}\}$, which contains all $(i,j)$ pairs where $(i,j)\in E$
@@
6:Return a classical vector $\boldsymbol{x}\leftarrow\boldsymbol{s}\cdot\sqrt{\boldsymbol{p}}$. $\sqrt{\boldsymbol{p}}$ means a element-wise square-root of the vector $\boldsymbol{p}$
4 Experiments on a real quantum computer
@@ 4.1 Experimental setup
To offer a comprehensive supplementary description of the computational process in this work, details of QLS have been added as follows:
  • 1.
    In this work, the initial solution $\tilde{\boldsymbol{x}}$ is always chosen as $\boldsymbol{0}$.
  • 2.
    The initial ansatz parameter $\boldsymbol{\theta}_{0}$ is determined as $\boldsymbol{\theta}_{0}={\arg\min_{\theta_{0}^{m}}}L(\boldsymbol{\theta}_{0}^{m})$, where $\boldsymbol{\theta}_{0}^{m}$ is generated by randomly sampling $m$ times from a uniform distribution over the interval $[0,2\pi)$.
  • 3.
    We utilized multiple learning rates $\lambda^{m}$, resulting in multiple $\boldsymbol{\theta}_{k+1}^{m}$ and $|\psi_{k+1}^{m}\rangle$. Then $L_{k+1}$ is chosen as $L_{k+1}=\min~{}\langle\psi_{k+1}^{m}|M|\psi_{k+1}^{m}\rangle$, and $|\psi_{k+1}\rangle$ is selected as $|\psi_{k+1}\rangle=\arg\min~{}\langle\psi_{k+1}^{m}|M|\psi_{k+1}^{m}\rangle$.
The experiments are conducted on a superconducting quantum computer, named “Wukong” which is fabricated by Origin Quantum. The detailed specification of Wukong is shown in A. No additional quantum error mitigation technique [52] is employed in experiments.
@@ 4.2 Numerical simulation of Poiseuille flow
where the superscript $+$ means the velocity in the next time step which is unknown, the subscript $j-1,j,j+1$ represents three adjacent grid points in the y-direction, $\Delta t$ is the discrete time step, and $\Delta y$ represents the grid size in the y-direction. Combining the discretized governing equations on all grid points in y-direction, we obtain a system of linear equations
@@
For a steady flow, it should hold that $\boldsymbol{x}^{+}=\boldsymbol{x}$, thus iteratively solving linear system Eq. (9) is equivalent to directly solving linear system as follows
$A\boldsymbol{x}=\boldsymbol{b},A=\begin{bmatrix}\beta&\alpha&0&0\\ \alpha&\beta&\ddots&0\\ 0&\ddots&\ddots&\alpha\\ 0&0&\alpha&\beta\\ \end{bmatrix},\quad\boldsymbol{b}=\gamma\begin{bmatrix}1\\ 1\\ \vdots\\ 1\\ \end{bmatrix},$ (10)
@@
In this work we set $p_{x}=-0.1,\mu=1,\rho=1,\Delta t=0.01$. The number of grids in the y-direction, denoted as $N_{y}$, is determined by the number of qubits used but always satisfies $N_{y}\Delta y=1$. For Iterative-QLS solver, we set the step size $\Delta=0.05$, loss function tolerance $\epsilon_{l}=10^{-3}$, convergence tolerance $\epsilon_{c}=10^{-4}$ and the number of shots for each circuit is $10^{4}$.
@@
A two-qubit experiment is conducted as a preliminary step to validate the feasibility of Iterative-QLS. From the results shown in Fig. 3(a), it can be seen that Iterative-QLS accurately solved the problem of 2D incompressible Poiseuille flow. The quantum results closely matched the analytical solutions with a relative error of less than 0.2%.
Figure 3: Experimental results of Poiseuille flow. (a) Quantitative comparison of velocity $u$ of 2D Poiseuille flow between results of Iterative-QLS solution $u_{q}$ (blue dots) on the with 2 qubits and analytical solutions $u_{a}$ (red line), where the relative error is defined as $|u_{q}/u_{a}-1|\times 10^{2}$. (b) Loss values change with increasing iteration steps. Each line segment is an independent VQLS execution, corresponding to an iteration step of Iterative-QLS. (c) Residual change with increasing iteration steps. The number marks the iteration step count, which shows the correspondence between (b) and (c).
@@
Fig. 3(b) illustrates the variation of the loss value during the training process. Each line segment represents an independent VQLS training process. Fig. 3(c) shows the change of loss during the iteration and Iterative-QLS reaches the convergence criterion after 4 iterations, with the residual decreasing to below $10^{-4}$. Here, the number marked in (b) corresponds to the iteration step count in (c). It can be seen that the residual value exhibits a clear decreasing trend, which confirms the successful execution of Iterative-QLS.
We also provide evidence of the scalability of Iterative-QLS, demonstrated with three-qubit and four-qubit experiments, shown in E.
@@ 4.3 Acoustic wave propagation simulation
In this section, linear acoustic wave propagation is simulated on NISQ devices. The acoustic wave propagation can be described with the linearized Euler equations as follows
@@
$S(t,x,y)=\epsilon p\ \exp^{-\alpha\frac{(x-x_{0})^{2}+(y-y_{0})^{2}}{b^{2}}}\sin(\omega t),$ (12)
where angle velocity $\omega$ and frequency f is related in $\omega=2\pi f$, and $(x_{0},y_{0})$ is the location of the sound source.
To accurately capture the sound waves, numerical schemes with low dissipation and low dispersion are required to discretize the equations. The dispersion relation preserving (DRP) scheme [53] is one of the most efficient schemes widely applied in computational aeroacoustics (CAA). In this work, the seven-point DRP scheme [54] is applied to spatial discretization. For time advance, the derivative is discretized with first order Euler implicit scheme to get higher advanced efficiency and more robust numerical stability. Furthermore, to avoid non-physical wave reflection near boundaries, the non-reflecting radiation boundary condition is imposed on all boundaries [55].
Applying the numerical methods above, the governing equations are discretized into the following form
@@
Fig. 5(a) shows a piece of convergence history of the subspace method during simulation. For each time step, the residual converges after 3 subspace iterations, showing the good convergence property of Iterative-QLS method. To further investigate the convergence performance of Iterative-QLS, a piece of convergence history of the iteration of VQLS during simulation is shown in Fig. 5(b). It is observed that, in each subspace iteration of every time step, the residual of Iterative-QLS drops quickly during iteration, showing the fast convergence of solving the linear system of equations with the subspace method.

### v3

@@ 3.3 Scaling quantum linear solver to arbitrary scale via the subspace method
To enhance scalability, we develop the subspace method inspired by the classical renowned Krylov subspace method and the generalized minimal residual (GMRES) method [49]. Instead of solving an $n$-dimensional linear system intractable large for near-term quantum computers, we iteratively build small spaces of lower dimension $m\ll n$ and find the approximate solution therein. More specifically, we consider the $m$-order Krylov subspace $\mathcal{K}_{m}$ spanned by basis $\left\{\boldsymbol{b},A\boldsymbol{b},A^{2}\boldsymbol{b},...,A^{m}\boldsymbol{b}\right\}$. This integration of Iterative-QLS and subspace method is implemented via a quantum-classical hybrid procedure. As shown in Fig. 1, the subspace is constructed on classical computers, and a subspace linear system is generated and solved by Iterative-QLS. The algorithm is presented in Alg. 2.
C provides theoretical details of the subspace method. We conduct convergence analysis and show that the algorithm can converge with a very loose error bound, i.e. $\epsilon<1$ for a single system solved by Iterative-QLS. Also, we show that the approximate solution $\hat{x}\in\mathcal{K}_{m}$ turns out to be a good estimation of $x$ when $m$ is large enough and with a restart scheme, as detailed in C.2.
Iterative-QLS is chosen as the internal solver for the subspace method primarily due to its high-precision features. Replacing Iterative-QLS with a single VQLS process may cause divergence in the subspace process, as each integration is not guaranteed to sufficiently shrink the residual, considering the error produced in a real quantum computer. Furthermore, Iterative-QLS can be replaced by a faster and more accurate quantum solver, which remains a subject for future works.
The size of the subspace can be determined by the capability of the internal QLS process. By providing a more powerful quantum computer, one can utilize a larger subspace size to accelerate the iteration process, an idea widely discussed in literature [50, 51, 52, 53, 54, 55]. Since QLS can solve $2^{n}$-dimensional linear system with $\operatorname{poly}(n)$ qubits and time complexity, it offers broad potential for quantum computing to accelerate a part of the fluid simulation procedure.
@@
1:Linear equations $A\boldsymbol{x}=\boldsymbol{b}$. Convergence criteria $\epsilon$. Subspace dimension $m$.
2:Give an initial solution $\tilde{\boldsymbol{x}}$.
3:Obtain the residual $\boldsymbol{r}=\boldsymbol{b}-A\tilde{\boldsymbol{x}}$ and $\beta=\|\boldsymbol{r}\|_{2}$.
4:while $\beta>\epsilon$ do
5:     Construct subspace linear system $H\boldsymbol{y}=\beta\boldsymbol{e}_{1}$ with coefficient matrix $V$ by $m$-order GMRES method.
6:     Solve the subspace linear system through Iterative-QLS.
7:     Update $\tilde{\boldsymbol{x}}=\tilde{\boldsymbol{x}}+V\boldsymbol{y}$.
8:     Obtain the residual $\boldsymbol{r}=\boldsymbol{b}-A\tilde{\boldsymbol{x}}$ and $\beta=\|\boldsymbol{r}\|_{2}$.
9:end while
@@ 3.4 Improving readout via sparse tomography
The concept of sparse sampling has been adopted in many domains which require fast data conversion from quantum to classical. $l_{\infty}$-norm tomography was the first proposed in constructing a deep convolutional neural network [48]. It offered an efficient method to generate such a sparse sample with $\mathcal{O}(\log N/\epsilon^{2})$ time complexity, where the error of each entry is less than $\epsilon$. It was further applied to quantum algorithms [56, 57, 58] that involve iterations where each iteration must convert quantum data to classical. In Ref. [56], the application of sparse tomography to various CFD problems is studied, showing that sparse tomography, acting as an “importance-sampling” scheme, is available for the quantum-enhanced finite volume method.
However, $l_{\infty}$-tomography requires quantum random access memory (QRAM) [59] to store the output and perform state preparation, as well as an additional qubit, resulting in extra costs for implementing the controlled version of the entire process. However, QRAM is widely believed to be unrealistic in the near term [60], thus $l_{\infty}$-tomography is impracticable in near-term quantum computers. To address these limitations, a new method called “sparse tomography” is introduced in this study to achieve the same sparse sampling task without utilizing QRAM or any additional qubits.
The major idea of sparse tomography is the exploitation of the shadow tomography [61, 62], which is a novel tool to measure $O(\log N)$ times to compute $N$ expectation values by exploiting a classical postprocessing procedure named “classical shadow”. In our algorithm, we construct the classical shadow by defining a computation graph to determine the sign of the quantum state. The algorithm’s procedure is shown in Alg. 3 and the complexity and proof can be found in D. Because the algorithm does not use QRAM and it only involves measurements on various measurement bases (refer to Ref. [62] for the concept about measurement bases), this procedure is efficient and completely capable for NISQ devices.
@@
1:A quantum algorithm that produces a $N$-dimensional real-valued state $|x\rangle$. Error criteria $\epsilon$.
2:Sample $|x\rangle$ with $36\log N/\epsilon^{2}$ copies to produce a sparse probability vector $\boldsymbol{p}$, and the position vector $\boldsymbol{v}$.
3:Construct graph $\mathcal{G}=(V,E)$ with $\log\epsilon/\log v(1-p)$ random connected spanning subgraphs, where each subgraph contains all nodes with least edges ($N-1$ edges). For each subgraph, prepare all sign-determination Hamiltonians $\{H_{X}^{i,j}\}$, which contains all $(i,j)$ pairs where $(i,j)\in E$.
@@
6:Return a classical vector $\boldsymbol{x}\leftarrow\boldsymbol{s}\cdot\sqrt{\boldsymbol{p}}$. $\sqrt{\boldsymbol{p}}$ means a element-wise square-root of the vector $\boldsymbol{p}$.
4 Experiments on a quantum computer
@@ 4.1 Experimental setup
We mainly utilize VQLS as the base solver for building Iterative-QLS. To offer a comprehensive supplementary description of the computational process in this work, we provide a tutorial of VQLS in B.2. In the VQLS process, the initial solution $\tilde{\boldsymbol{x}}$ is always chosen as $\boldsymbol{0}$. The initial ansatz parameter $\boldsymbol{\theta}_{0}$ is determined as $\boldsymbol{\theta}_{0}={\arg\min_{\theta_{0}^{m}}}L(\boldsymbol{\theta}_{0}^{m})$, where $\boldsymbol{\theta}_{0}^{m}$ is generated by randomly sampling $m$ times from a uniform distribution over the interval $[0,2\pi)$.
The experiments are conducted on a superconducting quantum computer, named “Wukong” which is fabricated by Origin Quantum. The detailed specification of Wukong is shown in A. No additional quantum error mitigation technique [27] is employed in the following experiments.
(a)(b)(c)1234init1234
Figure 3: Experimental results of Poiseuille flow. (a) Quantitative comparison of velocity $u$ of 2D Poiseuille flow among results of Iterative-QLS solution $u_{q}$ (blue square) on four points with two qubits, single execution of VQLS (red dots), and analytical solutions $u_{a}$ (black line). Relative error for Iterative-QLS is also plotted, defined as $|u_{q}/u_{a}-1|\times 10^{2}$. (b) Loss values change with increasing iteration steps. Each line segment is an independent VQLS training process, corresponding to an iteration step of Iterative-QLS. (c) Residual change with increasing iteration steps. The number marks the iteration step count, which shows the correspondence between (b) and (c).
@@ 4.2 Numerical simulation of Poiseuille flow
where the superscript $+$ means the velocity in the next time step which is unknown, the subscript $j-1$, $j$, and $j+1$ represent three adjacent grid points in the y-direction, $\Delta t$ is the discrete time step, and $\Delta y$ represents the grid size in the y-direction. Combining the discretized governing equations on all grid points in the y-direction, we obtain a system of linear equations
@@
For a steady flow, it should hold that $\boldsymbol{x}^{+}=\boldsymbol{x}$, thus iteratively solving linear system Eq. (11) is equivalent to directly solving linear system as follows:
$A\boldsymbol{x}=\boldsymbol{b},\quad A=\begin{bmatrix}\beta&\alpha&0&0\\ \alpha&\beta&\ddots&0\\ 0&\ddots&\ddots&\alpha\\ 0&0&\alpha&\beta\\ \end{bmatrix},\quad\boldsymbol{b}=\gamma\begin{bmatrix}1\\ 1\\ \vdots\\ 1\\ \end{bmatrix},$ (12)
@@
In this work, we set $p_{x}=-0.1$, $\mu=1$, $\rho=1$, and $\Delta t=0.01$. The number of grids in the y-direction, denoted as $N_{y}$, is determined by the number of qubits used but always satisfies $N_{y}\Delta y=1$. For Iterative-QLS, we set the loss function tolerance $\epsilon_{l}=10^{-3}$, convergence tolerance $\epsilon_{c}=10^{-4}$ and the number of shots for each circuit is $10^{4}$.
@@
A two-qubit experiment is conducted as a preliminary step to validate the feasibility of Iterative-QLS. From the results shown in Fig. 3(a), it can be seen that Iterative-QLS accurately solved the problem of 2D incompressible Poiseuille flow. The quantum results closely matched the analytical solutions with a relative error of less than 0.2%. We also plot the results produced by a single execution of VQLS. By comparing the results of Iterative-QLS and a single execution of VQLS, we conclude that Iterative-QLS can effectively suppress the error generated in VQLS.
@@
Fig. 3(b) illustrates the variation of the loss value during the training process. Each line segment represents an independent VQLS training process. Fig. 3(c) shows the change of loss during the iteration and Iterative-QLS reaches the convergence criteria after 4 iterations, with the residual decreasing to below $10^{-4}$. Here, the number marked in subfigure (b) corresponds to the iteration step count in subfigure (c). It can be seen that the residual value exhibits a clear decreasing trend, which confirms the successful execution of Iterative-QLS.
(a)(b)
Figure 4: Experimental results of Poiseuille flow with more qubits. (a) Experimental results for 3-qubit Poiseuille simulation on 8 points. (b) Experimental results for 4-qubit Poiseuille simulation on 16 points. We plot results of Iterative-QLS solution $u_{q}$ (blue square) on four points with two qubits, single execution of VQLS (red dots), and analytical solutions $u_{a}$ (black line). Relative error for Iterative-QLS is also plotted, defined as $|u_{q}/u_{a}-1|\times 10^{2}$.
(a)(b)
Figure 5: Convergence history of Poiseuille flow experiments. (a) Loss values change with increasing iteration steps. One can observe that the average loss increases with a growing number of qubits. (b) Residual change with increasing iteration steps. Note that two-qubit results shown in Fig. 4 are also plotted here as a comparison.
We also demonstrate 3-qubit and 4-qubit experiments, as shown in Fig. 4. The results also show qualitative consistency between the analytical solution and the experimental results. Compared to a single execution of VQLS, Iterative-QLS can indeed output results with higher precision. We observe that the performance decreases with an increasing number of qubits. To understand this phenomenon, we compare the convergence history from 2-qubit to 4-qubit in Fig. 5. The training procedure of VQLS, as demonstrated in subfigure (a), displays the original solution quality produced by each single VQLS, which clearly shows that more qubits will result in a decrease in the performance of each VQLS. Although Iterative-QLS manages to suppress the error of all cases, as shown in subfigure (b), the increasing error from single VQLS finally leads to a decreasing performance of the convergence.
Iterative-QLS utilizes iterations to suppress the error generated in VQLS. A similar counterpart is quantum error mitigation (QEM) [27], which also utilizes redundant circuits to achieve lower error on NISQ devices. Compared to QEM, our technique focuses more specifically on the linear solver problem, resulting in a much lower overhead. The number of circuits executed in two-qubit, three-qubit, and four-qubit experiments is less than 20, 60, and 100, respectively. QEM techniques like probabilistic error cancellation [63] or zero-noise extrapolation [28] should first introduce a heavy-cost error calibration circuit, then utilize thousands of circuit instances to obtain one result. Therefore, our technique is much more efficient than QEM in our experiments. Moreover, QEM only reduces circuit-level error, while our method can further reduce the inherent trainability issue of VQLS [64]. Finally, QEM does not conflict with our method, which could be an effective way to scale Iterative-QLS when the size of the quantum circuit grows larger in the future.
Last but not least, Iterative-QLS is not limited to the current version of VQLS and our implementation. Iterative-QLS could produce more accurate results by designing an ansatz with better expressibility [65, 66], or other types of quantum solvers [32, 34, 35, 37], along with a stronger quantum computer in the future.
@@ 4.3 Acoustic wave propagation simulation
In this section, linear acoustic wave propagation is simulated on NISQ devices. The acoustic wave propagation can be described with the linearized Euler equations as follows:
@@
$S(t,x,y)=\epsilon\ \exp^{-\alpha\frac{(x-x_{0})^{2}+(y-y_{0})^{2}}{b^{2}}}\sin(\omega t),$ (14)
where angle velocity $\omega$ and frequency f is related in $\omega=2\pi f$. $(x_{0},y_{0})$ is the location of the sound source and $\epsilon$ is the source amplitude.
To accurately capture the sound waves, numerical schemes with low dissipation and low dispersion are required to discretize the equations [67]. The dispersion relation preserving (DRP) scheme [68] is one of the most efficient schemes widely applied in computational aeroacoustics (CAA). The seven-point DRP scheme [69] is applied to spatial discretization in this work. The derivative is discretized with a first-order Euler implicit scheme for time advancement to achieve higher advanced efficiency and more robust numerical stability. Furthermore, to avoid non-physical wave reflection near boundaries, the non-reflecting radiation boundary condition is imposed on all boundaries [70].
Applying the numerical methods above, the governing equations are discretized into the following form:
@@
To explore the convergence property, we show the convergence history of the subspace method, including the external iteration by the subspace method and internal iteration by Iterative-QLS method. Fig. 7(a) shows a part of the convergence history of the external iterations during simulation. For each time step, the residual converges after 3 subspace iterations, showing the good convergence property of the subspace method. To further investigate the convergence performance of internal iterations, a piece of convergence history of the iteration of Iterative-QLS during simulation is shown in Fig. 7(b). It is observed that, in each subspace iteration of every time step, the residual of Iterative-QLS drops quickly during iteration, showing the efficient convergence of solving the linear system of equations with the subspace method.